# Reliably Detecting Model Failures in Deployment Without Labels

### v3

@@ A.1 Overview
A.2 Post-Deployment Deterioration (PDD) and Disagreement-based
Post-Deployment Deterioration (D-PDD)
@@ A.5 Proofs
We use the generalization bound for agnostic learning in Shalev-Shwartz and Ben-David [2014].
@@
We use the generalization bound for agnostic learning in Shalev-Shwartz and Ben-David [2014].
@@
We use the generalization bound for agnostic learning case Shalev-Shwartz and Ben-David [2014].
@@ B.1 Baselines Details
    Deep Kernel MMD (MMD-D, Liu et al. [2020]) The algorithm first learns a deep kernel by optimizing a criterion which yields the most powerful hypothesis test. With this learned kernel, permutation tests are run multiple times in order to determine a true positive rate for the algorithm. We interface the authors’ original source code with our repository and recycle their training procedures. Theirs can be found at https://github.com/fengliu90/DK-for-TST.
@@
    H-Divergence (H-Div, Zhao et al. [2022]) The algorithm fits Gaussian kernel density estimates for $\bm{P}_{x}$, $\bm{Q}_{x}$, and their uniform mixture $(\bm{P}_{x}+\bm{Q}_{x})/2$. Then, permutation tests are performed using the test statistic $H_{\ell}((\bm{P}_{x}+\bm{Q}_{x})/2)-\operatorname{min}\{H_{\ell}(\bm{Q}_{x}),H_{\ell}(\bm{P}_{x})\}$ where $H_{\ell}$ is the H-entropy with $\ell(x,a)$ the negative log likelihood of $x$ under distribution $a$ estimated by the Gaussian kernel density, in order to determine a true positive rate for the algorithm. This test statistic is an empirical estimate of the H-Min divergence. The choice of the particular H-divergence is a hyperparameter and is problem dependent, as well as the choice for how to generatively model the data distributions. The original paper further experimented with fitting Gaussian distributions as well as variational autoencoders (VAEs), both of which are not explored here. We interface the authors’ original source code with our repository. Theirs can be found at https://github.com/a7b23/H-Divergence/tree/main.
@@
    $f$-Divergences (KL-Div for KL-Divergence, JS-Div for Jensen-Shannon Divergence, Acuna et al. [2021]) $f$-divergence generalizes several notions of distances between probability distributions commonly used in machine learning. In this paper, we convert the Kullback-Leibler (KL) and the Jensen-Shannon (JS) divergences, particular cases of $f$-divergences, into permutation tests. More specifically, we first fit Gaussians on samples coming from $\bm{P}_{x}$ and $\bm{Q}_{x}$ using maximum likelihood. In the case of KL-divergence, the empirical KL-divergence is computed between the fitted Gaussians whereas for the JS-divergence, we fit an additional Gaussian on the mixture distribution $\bm{M}$ and leverage the identity:
@@
    Black Box Shift Detection (BBSD, Lipton et al. [2018]) involves estimating the changes in the distribution of target labels $p(y)$ between training and test data while assuming that the conditional distribution of features given labels $p(x|y)$ remains constant. This is achieved by using a black box model’s confusion matrix to identify discrepancies in the marginal label probabilities between the training and test distributions, allowing detection and correction of the shift. We borrow the experimental setup directly from Rabanser et al. [2019].
@@
    Relative Mahalanobis Distance (RMD, Rel-MH, Rel. Mahalanobis, Ren et al. [2021]) RMD modifies the traditional Mahalanobis Distance (MD) for out-of-distribution (OOD) detection by accounting for the influence of non-discriminative features. It subtracts the MD of a test sample to a background class-independent Gaussian from the MD to each class-specific Gaussian, effectively isolating discriminative features and improving OOD detection, especially for near-OOD tasks. We test for shift by performing a KS test directly on the distribution of the RMD confidence scored computed on $\bm{Q}_{x}$ and $\bm{P}_{x}$.
@@
    Classifier two-sample test (CTST, C2ST, a.k.a. Domain Classifier, DC, Lopez-Paz and Oquab [2016]) Following the procedure presribed by the original work, a binary domain classifier is trained to predict whether a sample came from $\bm{P}_{x}$ or $\bm{Q}_{x}$. Then, on a held-out mixture of data from $\bm{P}_{x}$ and $\bm{Q}_{x}$, we compute the domain classifier’s accuracy and compare its performance to random chance using a binomial test.
@@
    Deep Ensemble (Deep Ensemble, Ovadia et al. [2019]) An ensemble of neural networks are trained independently on the entire training dataset using random initialization. A KS test is then performed on the distribution of entropy values computed from each sample in $\bm{P}$ and $\bm{Q}$.
@@ B.2 Standard Benchmark Datasets
The UCI Heart Disease dataset (UCI) Janosi et al. [1989], Asuncion et al. [2007] includes 76 variables gathered from four distinct patient cohorts located in Cleveland, Hungary, Switzerland, and the VA Long Beach. To reduce the impact of missing data, we focus on nine out of the 14 most frequently used features: age, sex, chest pain type, resting blood pressure, serum cholesterol, fasting blood sugar, resting ECG results, maximum heart rate achieved, and exercise-induced angina. The objective is to predict heart disease diagnosis, measured on a scale from 0 to 4—where 0 denotes no disease and values 1 through 4 indicate increasing severity related to arterial narrowing. Using the proposed setup in Ginsberg et al. [2023], the task is binarized, distinguishing between patients with a normal angiographic diagnosis (label 0) and those with any abnormal diagnosis (label greater than 0). The Cleveland and Hungary datasets serve as the ID domain, while Switzerland and VA Long Beach datasets form the deteriorating OOD domain.
@@
The CIFAR-10 image dataset Krizhevsky et al. [2009] consists of 60,000 color images, each sized 32x32 pixels, divided into 10 different classes such as airplanes, cars, birds, and dogs. The dataset is split into 50,000 training images and 10,000 test images, with each class represented equally. Due to its manageable size and diversity of categories, CIFAR-10 is commonly used for testing and comparing the performance of image recognition models. The CIFAR-10.1 dataset Recht et al. [2018] is a test set designed to evaluate how well models trained on CIFAR-10 generalize to new, but similar, data. It consists of 2,000 images collected in a way that closely matches the original CIFAR-10 distribution, but from a separate data source to reduce the risk of overlap or memorization. CIFAR-10.1 was introduced to assess model robustness and identify potential overfitting to the original test set. Despite its similarity to CIFAR-10, many models perform slightly worse on CIFAR-10.1, highlighting the challenge of generalization in machine learning.
@@
Figure 5: (Left) Random samples from CIFAR-10.1. (Right) Random samples from CIFAR-10’s test set. The images above are borrowed from Recht et. al. “Do CIFAR-10 Classifiers Generalize to CIFAR-10?” [Recht et al., 2018].
@@
The Camelyon17 dataset is a challenging histopathology image classification dataset originally described by Bandi et al. [2018]. It comprises 327,680 color images (96×96) extracted from lymph node tissue slides, with binary labels indicating the presence of metastatic cancer in the central 32×32 region. Following the WILDS setup, we treat the hospitals from which data was collected as domains: hospitals 1, 2, and 3 are used as source domains for training, while hospital 5 serves as the target test domain. We use the WILDS framework Koh et al. [2021] to handle dataset download, preprocessing, and domain partitioning.
Figure 6: Samples from the Camelyon17 dataset, where hospitals 1, 2, 3 are treated as ID, while hospitals 4 and 5 are deterioratingly OOD. The image above is borrowed from Koh and Sagawa et. al. “WILDS: A benchmark of in-the-wild distribution shifts” [Koh et al., 2021].
@@ B.3 The GEneral Medicine INpatient Initiative (GEMINI) Dataset
The GEMINI study is a retrospective cohort study of adult patients and their clinical and administrative data[Verma et al., 2017, 2021]. This analysis used data from over 200,000 patients from the GEMINI database, spanning 7 different hospitals that participated in the GEMINI Study. Each patients information is processed into  900 features including but not limited to: (i) laboratory results and vital results collected up to 48-hours after admission, split into 6 hour intervals, (ii) patient demographic information: age, sex etc, (iii) Patient diagnosis using ICD-10-CA codes. Missing feature values are imputed based on simple averaging. The predictive task related to this data is 14-day mortality for patients based on these collected features. This dataset is temporal by nature as data splits are arranged per half-years starting from 2017, where shifts in patient features become even more pronounced during the COVID period of 2020-2021. Since our D3M base models trained on data before 2018 surprisingly do not underperform on later splits (Figure 2), we make the assessment that this shift is non-deteriorating.
@@ B.4 Configurations for $\operatorname{FE}_{\theta}$ and $\operatorname{VBLL}_{\theta}$
In the tabular setting, we chain affine layers of similar hidden dimension coupled with exponential linear unit (ELU) non-linearities [Clevert et al., 2015]. We use a standard dropout rate of $0.2$ across all experiments and employ skip connections between hidden layers. This architecture is used in all experiments with the UCI Heart Disease dataset as well as the GEMINI dataset.
@@
For CIFAR-10/10.1, we pass the input through an initial convolution and max-pooling, then through a sequence of same-dimension convolutions with batch normalization Ioffe and Szegedy [2015] and skip-connections. Another max-pool is applied, before the representation is flattened and forwarded through an affine layer to obtain a representation.
For Camelyon17 experiments, we employ pre-trained ResNets He et al. [2016] and either train from scratch, finetune them during the training of the $\operatorname{VBLL}_{\theta}$ layer, or freeze them while only training the VBLL layer.
@@
We borrow the implementation from Harrison et al. [2024] which can be readily coupled with the above neural feature extractors for end-to-end ELBO maximization. In our experiments, we employ the VBLL variant for discriminative classification vbll.DiscClassification, and parametrize the covariance matrix of the normal distribution of logits as a diagonal. As for VBLL’s hyperparameters, the prior scale and wishart scale hyperparameters are as described in the original manuscript, while we use the regularization factor to VBLL to be a factor of $n^{-1}$ where $n$ is the size of the training set. This factor controls the weight of the KL estimate during ELBO maximization at training.
@@ B.5 Sweeping and Hyperparameters
For all experiments, the number of posterior samples $K$ is set to $5000$, the size of the empirical distribution of maximum disagreement rates $|\Phi|$ is set to $1000$. For each experiment, once Train and Calibrate is completed, we deploy the model on $100$ independent samplings of the questionable test data. If model deterioration occurs (this is the case for the standard benchmark experiments as well as the GEMINI dataset temporal shift experiment but not the GEMINI dataset age shift experiment), we report the number of times D3M flagged these deteriorating samples out of $100$ as TPR. Finally, confidence statistics are aggregated and computed on TPRs reported from seeded, independent Train-Calibrate-Deploy D3M cycles. All optimization is done using AdamW Loshchilov and Hutter [2017]. Finally, we start seeded, independent, identical runs beginning at $\operatorname{seed}=57$ since it is our favorite prime number, and increase $\operatorname{seed}$ by +1+1 for each run.
@@ B.6 Additional D3M Results on Standard Benchmark
D3M is efficient when comparing to other disagreement-based detection and monitoring algorithms. However, the method is inherently noisier due to several levels of approximation to its idealized version (Algorithms 1 and 2). In particular, we found that D3M is sensitive to the choice of hyperparameters in order to achieve great performance. Therefore, the payment of computational cost is carried through prior to deployment in the sweeping itself, rather than during deployment as is done in Ginsberg et al. [2023]. We argue, however, that this is preferable in edge deployment scenarios, such as in hospitals or embedded systems, where real-time responsiveness and resource constraints are critical by front-loading the computational burden during the sweeping phase. One can imagine that for an AI terminal as part of a hospital computational infrastructure for instance, the ability to perform robust detection and monitoring with minimal overhead at deployment time significantly enhances reliability and usability in practice.
@@ C.4 Oversampling of the ideal restricted hypothesis space $\mathcal{H}_{p}$
In contrast, Detectron [Ginsberg et al., 2023] optimizes a joint loss term trading off between a training error minimization and a target unsupervised dataset’s error maximization (with respect to its pseudolabels). As this tradeoff is tunable, it’s possible to enforce stronger adherence to hypotheses within $\mathcal{H}_{p}$ at the heavy price of requiring the training set to be at arm’s length.

### v4

@@ 5 Conclusion and Limitations
Acknowledgments and Disclosure of Funding
RGK gratefully acknowledges support from the Canada Research Chairs Program (CRC-2022-00049), the Canada CIFAR AI Chairs Program. This research was funded in part by a NFRF Special Call Award (NFRFR-2022-00526). Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute (www.vectorinstitute.ai/partnerships/current-partners/). This study was supported by GEMINI, a research program based out of Unity Health Toronto. The development of GEMINI’s data platform has been supported by several funding partners, which can be found listed at https://geminimedicine.ca/partners/. All analyses, results, conclusions, and opinions presented in this paper are solely those of the listed authors and are independent from GEMINI’s funding sources. No endorsement by GEMINI’s funding sources is intended, nor should it be inferred. Amol Verma receives salary support from the Temerty Professorship of AI Research and Education in Medicine at the University of Toronto. Fahad Razak receives salary support from the Canada Research Chairs program. VN gratefully acknowledges the invaluable discussions and the many thoughtful tea sessions with Michael Cooper and Vahid Balazadeh.
Appendix
@@ A.1 Overview
A.2 Post-Deployment Deterioration (PDD) and Disagreement-based Post-Deployment Deterioration (D-PDD)
@@ A.5 Proofs
We use the generalization bound for agnostic learning in [63].
@@
We use the generalization bound for agnostic learning in [63].
@@
We use the generalization bound for agnostic learning case [63].
@@ B.1 Baselines Details
    Deep Kernel MMD (MMD-D, [13]) The algorithm first learns a deep kernel by optimizing a criterion which yields the most powerful hypothesis test. With this learned kernel, permutation tests are run multiple times in order to determine a true positive rate for the algorithm. We interface the authors’ original source code with our repository and recycle their training procedures. Theirs can be found at https://github.com/fengliu90/DK-for-TST.
@@
    H-Divergence (H-Div, [14]) The algorithm fits Gaussian kernel density estimates for $\bm{P}_{x}$, $\bm{Q}_{x}$, and their uniform mixture $(\bm{P}_{x}+\bm{Q}_{x})/2$. Then, permutation tests are performed using the test statistic $H_{\ell}((\bm{P}_{x}+\bm{Q}_{x})/2)-\operatorname{min}\{H_{\ell}(\bm{Q}_{x}),H_{\ell}(\bm{P}_{x})\}$ where $H_{\ell}$ is the H-entropy with $\ell(x,a)$ the negative log likelihood of $x$ under distribution $a$ estimated by the Gaussian kernel density, in order to determine a true positive rate for the algorithm. This test statistic is an empirical estimate of the H-Min divergence. The choice of the particular H-divergence is a hyperparameter and is problem dependent, as well as the choice for how to generatively model the data distributions. The original paper further experimented with fitting Gaussian distributions as well as variational autoencoders (VAEs), both of which are not explored here. We interface the authors’ original source code with our repository. Theirs can be found at https://github.com/a7b23/H-Divergence/tree/main.
@@
    $f$-Divergences (KL-Div for KL-Divergence, JS-Div for Jensen-Shannon Divergence, [61]) $f$-divergence generalizes several notions of distances between probability distributions commonly used in machine learning. In this paper, we convert the Kullback-Leibler (KL) and the Jensen-Shannon (JS) divergences, particular cases of $f$-divergences, into permutation tests. More specifically, we first fit Gaussians on samples coming from $\bm{P}_{x}$ and $\bm{Q}_{x}$ using maximum likelihood. In the case of KL-divergence, the empirical KL-divergence is computed between the fitted Gaussians whereas for the JS-divergence, we fit an additional Gaussian on the mixture distribution $\bm{M}$ and leverage the identity:
@@
    Black Box Shift Detection (BBSD, [38]) involves estimating the changes in the distribution of target labels $p(y)$ between training and test data while assuming that the conditional distribution of features given labels $p(x|y)$ remains constant. This is achieved by using a black box model’s confusion matrix to identify discrepancies in the marginal label probabilities between the training and test distributions, allowing detection and correction of the shift. We borrow the experimental setup directly from [1].
@@
    Relative Mahalanobis Distance (RMD, Rel-MH, Rel. Mahalanobis, [39]) RMD modifies the traditional Mahalanobis Distance (MD) for out-of-distribution (OOD) detection by accounting for the influence of non-discriminative features. It subtracts the MD of a test sample to a background class-independent Gaussian from the MD to each class-specific Gaussian, effectively isolating discriminative features and improving OOD detection, especially for near-OOD tasks. We test for shift by performing a KS test directly on the distribution of the RMD confidence scored computed on $\bm{Q}_{x}$ and $\bm{P}_{x}$.
@@
    Classifier two-sample test (CTST, C2ST, a.k.a. Domain Classifier, DC, [41]) Following the procedure presribed by the original work, a binary domain classifier is trained to predict whether a sample came from $\bm{P}_{x}$ or $\bm{Q}_{x}$. Then, on a held-out mixture of data from $\bm{P}_{x}$ and $\bm{Q}_{x}$, we compute the domain classifier’s accuracy and compare its performance to random chance using a binomial test.
@@
    Deep Ensemble (Deep Ensemble, [40]) An ensemble of neural networks are trained independently on the entire training dataset using random initialization. A KS test is then performed on the distribution of entropy values computed from each sample in $\bm{P}$ and $\bm{Q}$.
@@ B.2 Standard Benchmark Datasets
The UCI Heart Disease dataset (UCI) [21, 64] includes 76 variables gathered from four distinct patient cohorts located in Cleveland, Hungary, Switzerland, and the VA Long Beach. To reduce the impact of missing data, we focus on nine out of the 14 most frequently used features: age, sex, chest pain type, resting blood pressure, serum cholesterol, fasting blood sugar, resting ECG results, maximum heart rate achieved, and exercise-induced angina. The objective is to predict heart disease diagnosis, measured on a scale from 0 to 4—where 0 denotes no disease and values 1 through 4 indicate increasing severity related to arterial narrowing. Using the proposed setup in [19], the task is binarized, distinguishing between patients with a normal angiographic diagnosis (label 0) and those with any abnormal diagnosis (label greater than 0). The Cleveland and Hungary datasets serve as the ID domain, while Switzerland and VA Long Beach datasets form the deteriorating OOD domain.
@@
The CIFAR-10 image dataset [32] consists of 60,000 color images, each sized 32x32 pixels, divided into 10 different classes such as airplanes, cars, birds, and dogs. The dataset is split into 50,000 training images and 10,000 test images, with each class represented equally. Due to its manageable size and diversity of categories, CIFAR-10 is commonly used for testing and comparing the performance of image recognition models. The CIFAR-10.1 dataset [2] is a test set designed to evaluate how well models trained on CIFAR-10 generalize to new, but similar, data. It consists of 2,000 images collected in a way that closely matches the original CIFAR-10 distribution, but from a separate data source to reduce the risk of overlap or memorization. CIFAR-10.1 was introduced to assess model robustness and identify potential overfitting to the original test set. Despite its similarity to CIFAR-10, many models perform slightly worse on CIFAR-10.1, highlighting the challenge of generalization in machine learning.
@@
Figure 5: (Left) Random samples from CIFAR-10.1. (Right) Random samples from CIFAR-10’s test set. The images above are borrowed from Recht et. al. “Do CIFAR-10 Classifiers Generalize to CIFAR-10?” [2].
@@
The Camelyon17 dataset is a challenging histopathology image classification dataset originally described by [65]. It comprises 327,680 color images (96×96) extracted from lymph node tissue slides, with binary labels indicating the presence of metastatic cancer in the central 32×32 region. Following the WILDS setup, we treat the hospitals from which data was collected as domains: hospitals 1, 2, and 3 are used as source domains for training, while hospital 5 serves as the target test domain. We use the WILDS framework [4] to handle dataset download, preprocessing, and domain partitioning.
Figure 6: Samples from the Camelyon17 dataset, where hospitals 1, 2, 3 are treated as ID, while hospitals 4 and 5 are deterioratingly OOD. The image above is borrowed from Koh and Sagawa et. al. “WILDS: A benchmark of in-the-wild distribution shifts” [4].
@@ B.3 The GEneral Medicine INpatient Initiative (GEMINI) Dataset
The GEMINI study is a retrospective cohort study of adult patients and their clinical and administrative data[22, 23]. This analysis used data from over 200,000 patients from the GEMINI database, spanning 7 different hospitals that participated in the GEMINI Study. Each patients information is processed into  900 features including but not limited to: (i) laboratory results and vital results collected up to 48-hours after admission, split into 6 hour intervals, (ii) patient demographic information: age, sex etc, (iii) Patient diagnosis using ICD-10-CA codes. Missing feature values are imputed based on simple averaging. The predictive task related to this data is 14-day mortality for patients based on these collected features. This dataset is temporal by nature as data splits are arranged per half-years starting from 2017, where shifts in patient features become even more pronounced during the COVID period of 2020-2021. Since our D3M base models trained on data before 2018 surprisingly do not underperform on later splits (Figure 2), we make the assessment that this shift is non-deteriorating.
@@ B.4 Configurations for $\operatorname{FE}_{\theta}$ and $\operatorname{VBLL}_{\theta}$
In the tabular setting, we chain affine layers of similar hidden dimension coupled with exponential linear unit (ELU) non-linearities [66]. We use a standard dropout rate of $0.2$ across all experiments and employ skip connections between hidden layers. This architecture is used in all experiments with the UCI Heart Disease dataset as well as the GEMINI dataset.
@@
For CIFAR-10/10.1, we pass the input through an initial convolution and max-pooling, then through a sequence of same-dimension convolutions with batch normalization [67] and skip-connections. Another max-pool is applied, before the representation is flattened and forwarded through an affine layer to obtain a representation.
For Camelyon17 experiments, we employ pre-trained ResNets [37] and either train from scratch, finetune them during the training of the $\operatorname{VBLL}_{\theta}$ layer, or freeze them while only training the VBLL layer.
@@
We borrow the implementation from [25] which can be readily coupled with the above neural feature extractors for end-to-end ELBO maximization. In our experiments, we employ the VBLL variant for discriminative classification vbll.DiscClassification, and parametrize the covariance matrix of the normal distribution of logits as a diagonal. As for VBLL’s hyperparameters, the prior scale and wishart scale hyperparameters are as described in the original manuscript, while we use the regularization factor to VBLL to be a factor of $n^{-1}$ where $n$ is the size of the training set. This factor controls the weight of the KL estimate during ELBO maximization at training.
@@ B.5 Sweeping and Hyperparameters
For all experiments, the number of posterior samples $K$ is set to $5000$, the size of the empirical distribution of maximum disagreement rates $|\Phi|$ is set to $1000$. For each experiment, once Train and Calibrate is completed, we deploy the model on $100$ independent samplings of the questionable test data. If model deterioration occurs (this is the case for the standard benchmark experiments as well as the GEMINI dataset temporal shift experiment but not the GEMINI dataset age shift experiment), we report the number of times D3M flagged these deteriorating samples out of $100$ as TPR. Finally, confidence statistics are aggregated and computed on TPRs reported from seeded, independent Train-Calibrate-Deploy D3M cycles. All optimization is done using AdamW [68]. Finally, we start seeded, independent, identical runs beginning at $\operatorname{seed}=57$ since it is our favorite prime number, and increase $\operatorname{seed}$ by +1+1 for each run.
@@ B.6 Additional D3M Results on Standard Benchmark
D3M is efficient when comparing to other disagreement-based detection and monitoring algorithms. However, the method is inherently noisier due to several levels of approximation to its idealized version (Algorithms 1 and 2). In particular, we found that D3M is sensitive to the choice of hyperparameters in order to achieve great performance. Therefore, the payment of computational cost is carried through prior to deployment in the sweeping itself, rather than during deployment as is done in [19]. We argue, however, that this is preferable in edge deployment scenarios, such as in hospitals or embedded systems, where real-time responsiveness and resource constraints are critical by front-loading the computational burden during the sweeping phase. One can imagine that for an AI terminal as part of a hospital computational infrastructure for instance, the ability to perform robust detection and monitoring with minimal overhead at deployment time significantly enhances reliability and usability in practice.
@@ C.4 Oversampling of the ideal restricted hypothesis space $\mathcal{H}_{p}$
In contrast, Detectron [19] optimizes a joint loss term trading off between a training error minimization and a target unsupervised dataset’s error maximization (with respect to its pseudolabels). As this tradeoff is tunable, it’s possible to enforce stronger adherence to hypotheses within $\mathcal{H}_{p}$ at the heavy price of requiring the training set to be at arm’s length.